1370 hipass filter ripple in polA
lband wide has a filter bank installed after
the dewar. There are 9 filters switchable filters:
1 1370 hipass (ext Input unit. Use with
1280 hipass (ext Output unit. Use with filter 5)
Filters 2-8 can be or'd together (as long as the frequencies don't
Filters 1,9 are inline switchable before and after filters
2-8. They always need one of the 2-8 filters inserted to complete
the path (usually filter 5).
the lbfb command can be used to insert one or more filters
in the rf path.
- project a2703 reported ripples in polA of lbw on
- The period was about 10MHz.
- they used the interim correlator
- The ripple showed up in all 4 sbc of polA. This ruled out a
bad interim correlator board.
- They also used the 1100-1800 bandpass filter as well as the
1370 hipass filter.
- I looked at the ages data for 29oct12. This run used
- No ripple was seen.
- They used the wapps.
- They only used the 1100-1800 bandpass filter.
a2703 on/off ripple in spectral bandpass.
The first set of plots shows the ripple seen by a2703 on
276oct12 (.ps) (.pdf):
- Each page has the data for an on/off position switch pair.
these varied between 240 and 300 secs on.
- Black is pola, red is polB
- The top frame shows the spectral density function for the
first 25 MHz band (they had 2 offset in freq).
- the 2nd frame shows the median total power vs time for the 1
second samples of the on source
- the 3rd frame has the median total power vs time for the 1
second samples of the off source.
- The first on/off had no ripple.
- all the other on;off patterns show a large ripple in polA +/-
5% tsys up to < +/- 20% Tsys.
- PolA also showed a variation in the total power with time (up
to +/- 2% Tsys). PolB was relatively stable.
30oct12 testing the 1370 hipass filter looking at total pwr vs
On 30oct12 test data was taken, The setup was:
The first set of plots shows total
power vs time (.ps) (.pdf)
- lbw was used.
- the telescope was stationary.
- the interim correlator was used to take the data.
- 4 25 MHz bands were taken centered at: 1385,1400,1420,1440
- The problem was seen in all 4 bands.
- The total power in the 1385 MHz band was plotted vs time.
The 2nd set of plots uses
different rf switches for the 1370 hipass filter (.ps) (.pdf)
- Top: total power vs time using circular polarization.
- The circular is created after the dewar.
- At 200 seconds the 1370 filter was switched out. the
variation in polA stopped.
- Bottom: total power vs time using linear polarization (the
feed is natively linear).
- PolA is still the pol that is drifting.
- This means that the problem is after the hybrid (ie. it's
not in the dewar),.
- At 500, 600, and 760 secs datataking was stopped then
The these showed that the rf switches that 1370hp filter used seemed
to be the problem.
- total power vs time is plotted (linear polarization).
- the1370hp filter was in the path for all of these test
- Top: 1370hp filter using normal switches.
- Wiggling the filter input, output cables
- between green lines input cables were wiggled
- between blue lines output cables were wiggled.
- wiggling the cables made no difference.
- At 340 secs
- datataking was stopped (a source was drifting
- the command lbfb 1 , lbfb 1 5 was reissued.
- at 360 datataking was started and polA was drifting.
- 2nd: continue monitoring with normal switch positions
- 50 secs. stop.. a source had drifted through.
- issue commands, lbfb1 , lbfb 3, lbfb 1 5
- 80 secs: start datataking. polA is drifting again.
- 3rd: connect the 1370 hp filter to the switches normally used
by the 1290 hp filter.
- datataking was stopped, started 5 times.
- while stopped, different filters were switched in, out..
- the total power was stable. this points to the switched that
1370hp was using, not the filter itself.
- Bottom: 1370hp back to using its normal rf switches..
01nov12: install new rf switches for the 1370 highpass filter.
The rf switches that the 1370hp filter used were
The plots show the total
power vs time with the new switches (.ps) (.pdf)
- Top: total power vs time with 1370hp filter
- PolA drifts by about .5% for the 400 seconds. polB was
- at 500 secs
- datataking was stopped.
- the cmds lbfb 1 , lbfb 5, lbfb 1 5 , then datataking was
restarted. polA and B were stable.
- bottom check the stability of the other switches.
- datataking was setup to measure power at 1385 MHz with the
1100-1800 filter (filter 5)
- Before each scan (dashed blue line)
- filter 5 was swapped out
- filter N was inserted (N=1,9)
- filter 5 was switched back in
- This was looking to see if any of the other filters
interfered when they were switched in , then out.
- The dark blue line saw polA vary.
- This was filter 6 (1360->1470 bandpass) being switched
in and then out.
- a2703 saw ripples in their on/off band passes when the 1370
hipass filter was installed.
- Other users who did not use the 1370 hp filter did not have a
- Variations in total power always occurred in polA
(whether linear or circular pol was used).
- since the hybrid for circular is after the dewar, the
problem was not in the dewar.
- banging on the 1370hp input,output cable made no difference.
- switching the 1370hp filter to use the 1290hp filter rf
switches showed no variation in the power.
- Installing replacement rf switches for 1370hp made the total
power variation diminish to about .5% over 400 secs
- when switching in/out other filters, the 1360-1470 bandpass
filter seemed to affect the total power (even though it was not
- We may want to keep an eye on the lbw filter bank. The
switches may be getting old...(>15 years?).
- We should also make sure that the voltage from the switch
driver to the switches is correct.
processing: x101/121030/ chklbwbar.pro
lbw1370.pro lbwdbg.pro, 121101/lbwhpfilter.pro